OpenStudy (anonymous):

I need help with how to actually work this problem, not just an answer. f(x)=I 3x I +2x^2 even/odd/neither. I get odd, but when I plug it into an online algebra calculator, it spits out even. What am I doing wrong??

OpenStudy (anonymous):

the right and left side are mirror images of each other so its an even function

OpenStudy (anonymous):

yeah but how do you get that?

OpenStudy (anonymous):

I don't know how to plug in an absolute value into my calculator, is it possible?

OpenStudy (anonymous):

if its even then f(x)=f(-x) if its odd then f(x)=-f(-x)

OpenStudy (anonymous):

if u can prove f(x) is equal to f(-x) then it is an even function

OpenStudy (anonymous):

ok so I have to plug in a positive and a negative of the same number i.e. f(2) of f(-2) into the equation and either one will come up with an even number, correct?

OpenStudy (anonymous):

ok not necessarily even number but the SAME number

OpenStudy (anonymous):

yes then the function is an even function. u might want to try more than 1 number tho cuz it could just be coincidence

OpenStudy (anonymous):

yeah I've done a few and it's working out the same. The whole odd number was throwing me off. I thought the number had to be even, not the case.

OpenStudy (anonymous):

nope just the same but i think u got it now

OpenStudy (amistre64):



OpenStudy (amistre64):

you have one even function added to another; the result is an even function

OpenStudy (amistre64):

f(-x)= |-3x| +2(-x)^2 = |3x| +2(x)^2 = f(x) algebraically its even as well